Fermi National Laboratory
 

Operation Summary for October 18, 2010 - October 20, 2010

 

Monday Oct. 18

The day shift began with Operations monitoring the Tevatron (TeV) store 8179 (69.43E30) and the Recycler stash (343.9E10), with the Antiproton Source (Pbar) stacking (14.7E10), with the Linac, Booster, and the Main Injector (MI) operating normally, with NuMI off for target work, with the MiniBooNE and Meson FTBF CALICE T-978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Operations terminated store 8179 at 9:04 AM and then prepared the D0 collision hall for an access. After D0 came out of access, Operations put the TeV through a dry squeeze and wet squeeze. Operations put the TeV into shot setup and established store 8183 at 3:30 PM with an average initial luminosity (AIL) of 296.4E30. Operations reported issues with the Recycler Pelletron Electron Cooling (ECool) system suffering from full discharges, but operators resolved the problem before establishing store 8183.

During the evening shift, operators reported finding the TeV sector F4 pond water pump strainer completely clogged; the Duty Mechanic turned the pump off and set the system to run from a fire hydrant while repairs were made. Operations disabled the T-978 beam at 6 PM.

Tuesday Oct. 19

The midnight shift began with Operations monitoring store 8183 (36.20E30) and the stash (356.7E10), with Pbar stacking (29.4E10), and with MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8183 at 12:14 AM, put the TeV into shot setup, and established store 8184 at 1:40 AM with an AIL of 293E30. Operations established beam to T-978 at 6:25 AM.

During the day shift, Operations terminated store 8184 at 3:13 PM and put the TeV into shot setup. Operations reported several full discharges of the Pelletron; Pelletron personnel determined that a tank access would be needed after the TeV shot. Operations reported some downtime when the Booster Gradient Magnet Power Supply (GMPS) stopped regulating. Booster personnel changed out the regulator system and got GMPS running again.

During the evening shift, Operations reported that shot setup was brought to a screeching halt when the F0 vacuum system valves closed in sectors E4, F0, F1, A150, P150, and pulled the TeV abort. Vacuum technicians accessed F-Sector to repair some vacuum valve solenoids. Since the F-Sector access halted beam to all users, Safety personnel came in and replaced some beam detectors in Linac. Operations had the TeV ready to run by 10:49 PM, but MI beam was being held off by the MI-30 BLM system. Operators discovered the MI-30 rack 115 was without power, which was caused by a failed BLM crate power supply; Instrumentation techs came in and began repairs. Operations put the TeV through a dry squeeze during the MI BLM troubleshooting.

Wednesday Oct. 20

The midnight shift began with MI off due to a BLM problem, with Operations monitoring the stash (346.2E10) and the Pbar stack (10.4E10), and with MiniBooNE taking beam.

During the midnight shift, Instrumentation techs completed their repairs of MI-30 by 1:17 AM. Operations completed TeV orbit checks and had the TeV in shot setup by 1:52 AM, but that was delayed by an intermittent Booster RF problem. Operations established store 8188 at 3:17 AM with an AIL of 305.2E30. Operations had called in Booster RF techs to help with the problems encountered during shot setup, but the issue cleared up on its own shortly after the techs arrived. Operations prepared the Pelletron for access. Mechanical Support techs arrived at 4 AM to begin pumping the coolant out of the Pelletron tank.

The Plans for Wednesday

The plans for today are to monitor the store during the Pelletron repairs. The Run Coordinator will allow MiniBooNE personnel to access their hall, Pbar personnel to access their PreVault and Target areas for necessary repairs, and Booster will also access to repair some RF stations.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/20/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory